In the deck there are 4 sevens, 4 eights, 4 nines, 4 tens, 4 jacks, 4 queens, 4 kings, and 4 aces. What is the probability of pulling out one card and its a queen or a ace?
Answer by checkley75(3666)

8*4=32 total cards.
(4+4)/32=8/32=1/4=.25=25% chance of drawing a queen or an ace.